A farmer raises pigs & chickens. He has 220 animals with a total of 700 legs. How many pigs does the farmer have? Must show how. Thank you!
Answer by gsmani_iyer(178)


Let the No. of pigs = x
Let the no. of chicken = y
So, x + y = 220 ............... (1) (No. of animals)
4x + 2y = 700 .............. (2) (No. of total legs)
Eq.(1)*2, we get,
2x + 2y = 440 .............. (3)
Eq.(2) - (3), we get,
2x = 340
So x = 340/2 = 170
By replacing the value of x in eq.(1), we get
170 + y = 220
y = 220 - 170 = 50
So, x = 170; y = 50; which means 170 pigs and 50 chicken. Answer
I hope this is clear to you.


gsm